There are 300 purely competitive farms in the local dairy market. Of the 300 dairy farms, 298 have a cost structure that generates profits of $24 for every $300 invested.

Instructions: Enter your answers as whole numbers.

a. What is the percentage rate of return for these 298 dairies? percent.

b. The other two dairies have a cost structure that generates profits of $22 for every $200 invested. What is their percentage rate of return? percent.

c. Assuming that the normal rate of profit in the economy is 10 percent, and firms cannot copy each other's technology, will there be entry or exit? .

Will the change in the number of firms affect the two that earn $22 for every $200 invested? .

What will be the rate of return earned by most firms in the industry in long-run equilibrium? percent.

If firms can copy each other’s technology, what will be the rate of return eventually earned by all firms? percent.
1
Expert's answer
2020-04-27T07:59:29-0400

a)profit=$24

invested=$300

"\text{percentage rate of return}=(24/300)*100" =8%

b)profit=$22

invested=$200

"\text{percentage rate of return}=(22/200)*100%" =11%

c) rate= 10%

The other two firms will exit the market because the rate of returns would reduce.

If the number of firms reduce the effect would have positive effect to the 2 firms because there would be more market and vice versa.

If the firms copy each others technology the returns would be same ;

return=( 8%+11%)/2=9.5%